A rigid tank of volume 24 liters contains 0.2 kg of H2O originally at 2 bar. Find 

specific volume and initial temperature. What are the phases present (solid, 

liquid, vapor?). The tank was heated until its pressure reached 20 bar. Find work 

and heat exchanged during this process as well as final temperature


1
Expert's answer
2021-07-09T08:32:22-0400

Solution.

"V=24\sdot10^{-3}m^3;"

"m=0.2kg;"

"P_1=2\sdot10^5Pa;"

"P_2=20\sdot10^5Pa;"

"\dfrac{V}{m}=\dfrac{24\sdot10^{-3}}{0.2}=0.12m^3/kg;"

"M=18\sdot10^{-3}kg/mol;"

"PV=\dfrac{m}{M}RT\implies T_1=\dfrac{PVM}{mR};"

"T_1=\dfrac{2\sdot10^{5}\sdot24\sdot10^{-3}\sdot18\sdot10^{-3}}{0.2\sdot8.31}=52K;"

"\dfrac{P_1}{T_1}=\dfrac{P_2}{T_2};"

"T_2=\dfrac{T_1P_2}{P_1}=\dfrac{520\sdot20\sdot10^5}{2\sdot10^5}=520K;"

"W=0;"

"Q=c_Vm(T_2-T_1);"

"c_V=\dfrac{i}{2} R"


"Q=20.8\sdot0.2\sdot468=1947J;"

Answer: "\dfrac{V}{m}=0.12m^3/kg;"

"T_1=52K;"

"W=0;"

"Q=1947J."